OpenStudy (anonymous):

Which pair of fractions shows 7/9 and 5/12 correctly converted with their least common denominator?

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

They want both fractions to have the same denominators

OpenStudy (danjs):

to do this, take multiples of 12, until you get a number that 9 will divide evenly into

OpenStudy (anonymous):

Is it the same thing as last time?

OpenStudy (danjs):

sort of, you want to write both those fractions with the same denominators

OpenStudy (danjs):

the way i do it is, take multiples of the higher number, until you get a number that the smaller one will divide evenly into

OpenStudy (danjs):

12 * 2 = 24 9 will not divide evenly into 24

OpenStudy (anonymous):

I think it is c.

OpenStudy (danjs):

12*3 = 36 9 does divide into 36 evenly

OpenStudy (danjs):

36 is the denominators, now you need to multiply each fraction top and bottom by a number to make the denominators 36

OpenStudy (danjs):

7/9 * 4/4 = 28 / 36

OpenStudy (danjs):

5/12 * 3/3 = 15 / 36

OpenStudy (danjs):

see what i did there?

OpenStudy (anonymous):

34/36 and 29/36

OpenStudy (danjs):

no 28/36 and 15/36

OpenStudy (anonymous):

Is that the answer?

OpenStudy (anonymous):

K

OpenStudy (danjs):

yes, 36 is the least common denominator between 9 and 12

OpenStudy (anonymous):

That is right!!!!

OpenStudy (danjs):

multiplied the first fraction by 4/4 and the second by 3/3, then both will be over 36

OpenStudy (danjs):

it doesnt change the fraction at all 4/4 and 3/3 are just 1